# NeuralSurv: Deep Survival Analysis with Bayesian Uncertainty Quantification

### v1

@@ F.1 Approximating the Log Likelihood distribution
$\lambda^{\breve{\mathbb{Q}}}_{i}(t\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}):=\frac{t^{\rho-1}}{Z(t,\mathbf{x})}\phi^{(\ell)}\sigma(|g(t,\mathbf{x}_{i};\boldsymbol{\theta}^{(\ell)})|)\exp\left(-\frac{g(t,\mathbf{x}_{i};\boldsymbol{\theta}^{(\ell)})+|g(t,\mathbf{x}_{i};\boldsymbol{\theta}^{(\ell)})|}{2}\right).$ (33)
@@ F.2 EM Algorithm for MAP Estimation
    Latent Variables Update. Given the current estimates $(\boldsymbol{\theta},\phi)^{(\ell)}$, update $\breve{q}_{\boldsymbol{\omega}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}}(\boldsymbol{\omega}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)})$ and ˘$\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}}(\boldsymbol{\Psi}\mid\boldsymbol{\theta^{(\ell)}},\phi^{(\ell)})$ according to (31) and  (33), respectively.
@@
    $Q((\boldsymbol{\theta},\phi)|(\boldsymbol{\theta},\phi)^{(\ell)})=\\ \mathbb{E}_{\boldsymbol{\omega}\sim\breve{q}_{\boldsymbol{\omega}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}},\boldsymbol{\Psi}\sim\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}}}\left[\log\left(p(\mathcal{D}\mid\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi,\boldsymbol{\omega},\boldsymbol{\Psi})\>p_{\boldsymbol{\omega}}(\boldsymbol{\omega})\>\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}\mid\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})\right)\right]\\ +\log p_{\boldsymbol{\theta}}(\boldsymbol{\theta})+\log p_{\phi}(\phi).$ (34)
@@
7:         Update: $\left\{\breve{c}_{i}^{(\ell)}\right\}_{i=1}^{N}$ given $\boldsymbol{\theta}^{(\ell)}$ following (32).
8:     Update ˘$\breve{\mathbb{Q}}^{(\ell)}_{\boldsymbol{\Psi}}$:
9:         Update: $\left\{\lambda^{\breve{\mathbb{Q}},(\ell)}_{i}(\cdot)\right\}_{i=1}^{N}$ given $\boldsymbol{\theta}^{(\ell)}$ and $\phi^{(\ell)}$ following (40).
10:     E-step: Evaluate the Q-function $Q\bigl{(}(\boldsymbol{\theta},\phi)\mid(\boldsymbol{\theta},\phi)^{(\ell)}\bigr{)}$ given $\left\{\breve{c}_{i}^{(\ell)},\lambda^{\breve{\mathbb{Q}},(\ell)}_{i}(\cdot)\right\}_{i=1}^{N}$, $\boldsymbol{\theta}^{\ell}$ and $\phi^{\ell}$ following (34)
@@
The optimal distributions which minimize the KL divergence can now be plugged in the ELBO of (30) to obtain the closest lower bound to the log-likelihood. We now recast the MAP optimization problem in term of this lower bound. Specifically, define the following Q-function
@@
Next, by substituting the expression for the augmented likelihood in (13), for the priors $p_{\boldsymbol{\theta}}(\boldsymbol{\theta})$ in  (4) and $p_{\phi}(\phi)$ in  (5) and for the Radon–Nikodym derivative of $\mathbb{P}_{\boldsymbol{\Psi}\mid\phi}$ with respect to $\mathbb{P}_{\boldsymbol{\Psi},*}$ from (28), we obtain
@@
$-\frac{1}{2}\boldsymbol{\theta}^{\top}\boldsymbol{\theta}+\log(\phi)(\alpha_{0}-1)-\phi\beta_{0}+\text{const.}$.
We apply Campbell’s theorem (see Theorem C.3), we substitute the expression for the baseline hazard $\lambda_{i}(\cdot;\phi)$ from (11) and we substitute the expectation using the optimal variational distribution of $\omega_{i}$ from (31), to obtain
@@
$-\phi\left(\beta_{0}+\sum_{i=1}^{N}\int_{0}^{y_{i}}\frac{t^{\rho-1}}{Z(t,\mathbf{x}_{i})}\mathrm{d}t\right)+\text{const.},$. ,
where $\lambda^{\breve{\mathbb{Q}}}_{i}(t\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)})$ is shown in (33).
@@ G.1 Optimal Update for $\boldsymbol{\omega}$
Using standard mean-field variational inference techniques (see, e.g., Chapter 10.1 of [3]), the optimal update for the latent variables $\boldsymbol{\omega}$ is obtained by computing the expectation of the joint log-density with respect to the other variational factors. In particular, we have
@@
Finally, by applying the identity in (20), we deduce that the optimal variational distribution factorizes as
@@
$\tilde{c}^{(k)}_{i}=\delta_{i}\>\tilde{s}_{i}^{(k-1)}(y_{i})$ (35)
@@
$\mathbb{E}_{\omega_{i}\sim q^{(k)}_{\omega_{i}}}[\omega_{i}]=\frac{1}{2\tilde{c}^{(k)}_{i}}\text{tanh}\left(\frac{\tilde{c}^{(k)}_{i}}{2}\right)$ (36)
@@ G.2 Optimal Update for $\boldsymbol{\Psi}$
Using standard mean-field variational inference techniques (see, e.g., Chapter 10.1 of [3]), we obtain the optimal Radon-Nykodim derivative $\frac{\mathrm{d}\mathbb{Q}_{\boldsymbol{\Psi}}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}$ by taking the expectation of the joint log-density with respect to the other variational factors. In particular, we have
$\log\frac{\mathrm{d}\mathbb{Q}_{\boldsymbol{\Psi}}^{(k)}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})=\mathbb{E}_{\phi\sim q_{\phi}^{(k-1)},\boldsymbol{\theta}\sim q_{\boldsymbol{\theta}}^{(k-1)},\boldsymbol{\omega}\sim q_{\boldsymbol{\omega}}^{(k)}}\left[\log p(\mathcal{D}\mid\phi,g^{\text{lin}}(\cdot;\boldsymbol{\theta}),\boldsymbol{\omega},\boldsymbol{\Psi})\right]\\ +\mathbb{E}_{\phi\sim q_{\phi}^{(k-1)}}\left[\log\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}|\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})\right]+\text{const.},$ (37)
where the constant term absorbs all terms irrelevant to the optimisation. Using the augmented likelihood factorization in (14), the expression in (37) decomposes as
@@
Next, by substituting the augmented likelihood from (13) and the Radon–Nikodym derivative of $\mathbb{P}_{\boldsymbol{\Psi}|\phi}$ with respect to $\mathbb{P}_{\boldsymbol{\Psi},*}$ from (28), we arrive at the unnormalised form
$\log\frac{\mathrm{d}\mathbb{Q}_{\boldsymbol{\Psi}}^{(k)}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})=\sum_{i=1}^{N}\sum_{(t,\omega)_{j}\in\Psi_{i}}\mathbb{E}_{\boldsymbol{\theta}\sim q_{\boldsymbol{\theta}}^{(k-1)}}\left[f(\omega_{j},-g^{\text{lin}}(t_{j},\mathbf{x}_{i};\boldsymbol{\theta}))\right]\\ +\sum_{i=1}^{N}\sum_{(t,\omega)_{j}\in\Psi_{i}}\mathbb{E}_{\phi\sim q_{\phi}^{(k-1)}}\left[\log\left(\frac{\phi}{Z(t_{j},\mathbf{x}_{i})}\right)\right]+\text{const.}$ (38)
Plugging in the definition of $f(\cdot,\cdot)$ from (8) simplifies (38) to
@@
$h^{\mathbb{Q},(k)}_{i}(t,\omega)$ $:=\exp\left(-\frac{\left(\tilde{s}^{(k-1)}_{i}(t)\right)^{2}}{2}\omega\right)\,\cosh\left(\frac{\tilde{s}^{(k-1)}_{i}(t)}{2}\right),$
$\gamma_{i}^{\mathbb{Q},(k)}(t)$ $:=\frac{1}{Z(t,\mathbf{x}_{i})}\sigma(\tilde{s}^{(k-1)}_{i}(t))\exp\left(-\frac{\tilde{m}^{(k-1)}_{i}(t)+\tilde{s}^{(k-1)}_{i}(t)}{2}+\mathbb{E}_{\phi\sim q^{(k-1)}_{\phi}}[\log\phi]\right),$
$\lambda^{\mathbb{Q},(k)}_{i}(t,\omega)$ $:=\gamma_{i}^{\mathbb{Q},(k)}(t)h^{\mathbb{Q},(k)}_{i}(t,\omega)\lambda_{*,i}(t,\omega),$
where $\lambda_{*,i}(t,\omega)$ is the intensity defined in (26). Furthermore, we define for convenience,
$\lambda_{i}^{\mathbb{Q},(k)}(t):=t^{\rho-1}\gamma_{i}^{\mathbb{Q},(k)}(t).$ (39)
Notice that by using expression (20), the function $\lambda^{\mathbb{Q},(k)}_{i}(t,\omega)$ can be written as
$\lambda^{\mathbb{Q},(k)}_{i}(t,\omega)=\lambda^{\mathbb{Q},(k)}_{i}(t)\>p_{\text{PG}}\left(\omega\mid 1,\tilde{s}^{(k-1)}_{i}(t)\right).$ (40)
@@
Notice that the products $\gamma_{i}^{\mathbb{Q},(k)}(t_{j})h^{\mathbb{Q},(k)}_{i}(t_{j},\omega_{j})$ are all strictly positive  [^1] See Lemma N.1 for a proof of the strict positivity of the normalizing constant $Z(t,\mathbf{x}_{i})$., hence $\frac{\mathrm{d}\mathbb{Q}_{\boldsymbol{\Psi}}^{(k)}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}$ is also strictly positive. Under suitable regularity conditions on $g$, one can show that $h^{\mathbb{Q},(k)}_{i}(t,\omega)$ and $\gamma_{i}^{\mathbb{Q},(k)}(t)$ satisfy the integrability criteria of (27), so that $\mathbb{Q}_{\boldsymbol{\Psi}}^{(k)}$ is the probability measure under which each $\Psi_{i}$ $(i=1,\ldots,N)$ is a marked Poisson Process on $[0,y_{i}]\times\mathbb{R}_{+}$ with intensity function $\lambda_{i}^{\mathbb{Q},(k)}(t,\omega)$.
@@ G.3 Optimal Update for $\phi$
Using standard mean-field variational inference techniques (see, e.g., Chapter 10.1 of [3]), the optimal variational factor for the parameter $\phi$ is obtained by computing the expectation of the joint log-density with respect to the other variational factors. In particular, we have
@@
Next, by substituting the expression for the augmented likelihood from (13), the Radon–Nikodym derivative of $\mathbb{P}_{\boldsymbol{\Psi}|\phi}$ with respect to $\mathbb{P}_{\boldsymbol{\Psi},*}$ from (28), and the prior of $\phi$ from (5), we obtain,
@@
where $\lambda^{\mathbb{Q},(k)}_{i}(t)$ is shown in (39). We deduce that
$q^{(k)}_{\phi}(\phi)=\text{Gamma}(\tilde{\alpha}^{(k)},\tilde{\beta}),$
@@
$\tilde{\alpha}^{(k)}=\alpha_{0}+\sum_{i=1}^{N}\left(\delta_{i}+\int_{0}^{y_{i}}\lambda^{\mathbb{Q},(k)}_{i}(t)\mathrm{d}t\right),\quad\tilde{\beta}=\beta_{0}+\sum_{i=1}^{N}\int_{0}^{y_{i}}\frac{t^{\rho-1}}{Z(t,\mathbf{x}_{i})}\mathrm{d}t.$ (41)
@@
$\begin{split}&\mathbb{E}_{\phi\sim q^{(k)}_{\phi}}[\log\phi]=\psi\left(\tilde{\alpha}^{(k)}\right)-\log(\tilde{\beta}),\end{split}$ (42)

### v2

@@ F.1 Approximating the Log Likelihood distribution
$\lambda^{\breve{\mathbb{Q}}}_{i}(t\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}):=\frac{t^{\rho-1}}{Z(t,\mathbf{x})}\phi^{(\ell)}\sigma(|g(t,\mathbf{x}_{i};\boldsymbol{\theta}^{(\ell)})|)\exp\left(-\frac{g(t,\mathbf{x}_{i};\boldsymbol{\theta}^{(\ell)})+|g(t,\mathbf{x}_{i};\boldsymbol{\theta}^{(\ell)})|}{2}\right).$ (A15)
@@ F.2 EM Algorithm for MAP Estimation
    Latent Variables Update. Given the current estimates $(\boldsymbol{\theta},\phi)^{(\ell)}$, update $\breve{q}_{\boldsymbol{\omega}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}}(\boldsymbol{\omega}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)})$ and $\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}}(\boldsymbol{\Psi}\mid\boldsymbol{\theta^{(\ell)}},\phi^{(\ell)})$ according to (A13) and  (A15), respectively.
@@
    $Q((\boldsymbol{\theta},\phi)|(\boldsymbol{\theta},\phi)^{(\ell)})=\\ \mathbb{E}_{\boldsymbol{\omega}\sim\breve{q}_{\boldsymbol{\omega}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}},\boldsymbol{\Psi}\sim\breve{\mathbb{Q}}_{\boldsymbol{\Psi}\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)}}}\left[\log\left(p(\mathcal{D}\mid\mathbf{X},g(\cdot;\boldsymbol{\theta}),\phi,\boldsymbol{\omega},\boldsymbol{\Psi})\>p_{\boldsymbol{\omega}}(\boldsymbol{\omega})\>\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}\mid\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})\right)\right]\\ +\log p_{\boldsymbol{\theta}}(\boldsymbol{\theta})+\log p_{\phi}(\phi).$ (A16)
@@
7:    Update: $\left\{\breve{c}_{i}^{(\ell)}\right\}_{i=1}^{N}$ given $\boldsymbol{\theta}^{(\ell)}$ following (A14).
8:  Update $\breve{\mathbb{Q}}^{(\ell)}_{\boldsymbol{\Psi}}$:
9:    Update: $\left\{\lambda^{\breve{\mathbb{Q}},(\ell)}_{i}(\cdot)\right\}_{i=1}^{N}$ given $\boldsymbol{\theta}^{(\ell)}$ and $\phi^{(\ell)}$ following (A22).
10:  E-step: Evaluate the Q-function $Q\bigl((\boldsymbol{\theta},\phi)\mid(\boldsymbol{\theta},\phi)^{(\ell)}\bigr)$ given $\left\{\breve{c}_{i}^{(\ell)},\lambda^{\breve{\mathbb{Q}},(\ell)}_{i}(\cdot)\right\}_{i=1}^{N}$, $\boldsymbol{\theta}^{\ell}$ and $\phi^{\ell}$ following (A16)
@@
The optimal distributions which minimize the KL divergence can now be plugged in the ELBO of (A12) to obtain the closest lower bound to the log-likelihood. We now recast the MAP optimization problem in term of this lower bound. Specifically, define the following Q-function
@@
Next, by substituting the expression for the augmented likelihood in (14), for the priors $p_{\boldsymbol{\theta}}(\boldsymbol{\theta})$ in  (5) and $p_{\phi}(\phi)$ in  (2) and for the Radon–Nikodym derivative of $\mathbb{P}_{\boldsymbol{\Psi}\mid\phi}$ with respect to $\mathbb{P}_{\boldsymbol{\Psi},*}$ from (A10), we obtain
@@
$-\frac{1}{2}\boldsymbol{\theta}^{\top}\boldsymbol{\theta}+\log(\phi)(\alpha_{0}-1)-\phi\beta_{0}+\text{const.}$
We apply Campbell’s theorem (see Theorem C.3), we substitute the expression for the baseline hazard $\lambda_{i}(\cdot;\phi)$ from (12) and we substitute the expectation using the optimal variational distribution of $\omega_{i}$ from (A13), to obtain
@@
$-\phi\left(\beta_{0}+\sum_{i=1}^{N}\int_{0}^{y_{i}}\frac{t^{\rho-1}}{Z(t,\mathbf{x}_{i})}\mathrm{d}t\right)+\text{const.},$
where $\lambda^{\breve{\mathbb{Q}}}_{i}(t\mid\boldsymbol{\theta}^{(\ell)},\phi^{(\ell)})$ is shown in (A15).
@@ G.1 Optimal Update for $\boldsymbol{\omega}$
Using standard mean-field variational inference techniques (see, e.g., Chapter 10.1 of [bishop_pattern_recognition]), the optimal update for the latent variables $\boldsymbol{\omega}$ is obtained by computing the expectation of the joint log-density with respect to the other variational factors. In particular, we have
@@
Finally, by applying the identity in (A2), we deduce that the optimal variational distribution factorizes as
@@
$\tilde{c}^{(k)}_{i}=\delta_{i}\>\tilde{s}_{i}^{(k-1)}(y_{i})$ (A17)
@@
$\mathbb{E}_{\omega_{i}\sim q^{(k)}_{\omega_{i}}}[\omega_{i}]=\frac{1}{2\tilde{c}^{(k)}_{i}}\text{tanh}\left(\frac{\tilde{c}^{(k)}_{i}}{2}\right)$ (A18)
@@ G.2 Optimal Update for $\boldsymbol{\Psi}$
Using standard mean-field variational inference techniques (see, e.g., Chapter 10.1 of [bishop_pattern_recognition]), we obtain the optimal Radon-Nykodim derivative $\frac{\mathrm{d}\mathbb{Q}_{\boldsymbol{\Psi}}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}$ by taking the expectation of the joint log-density with respect to the other variational factors. In particular, we have
$\log\frac{\mathrm{d}\mathbb{Q}_{\boldsymbol{\Psi}}^{(k)}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})=\mathbb{E}_{\phi\sim q_{\phi}^{(k-1)},\boldsymbol{\theta}\sim q_{\boldsymbol{\theta}}^{(k-1)},\boldsymbol{\omega}\sim q_{\boldsymbol{\omega}}^{(k)}}\left[\log p(\mathcal{D}\mid\phi,g^{\text{lin}}(\cdot;\boldsymbol{\theta}),\boldsymbol{\omega},\boldsymbol{\Psi})\right]\\ +\mathbb{E}_{\phi\sim q_{\phi}^{(k-1)}}\left[\log\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}|\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})\right]+\text{const.},$ (A19)
where the constant term absorbs all terms irrelevant to the optimisation. Using the augmented likelihood factorization in (15), the expression in (A19) decomposes as
@@
Next, by substituting the augmented likelihood from (14) and the Radon–Nikodym derivative of $\mathbb{P}_{\boldsymbol{\Psi}|\phi}$ with respect to $\mathbb{P}_{\boldsymbol{\Psi},*}$ from (A10), we arrive at the unnormalised form
$\log\frac{\mathrm{d}\mathbb{Q}_{\boldsymbol{\Psi}}^{(k)}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}(\boldsymbol{\Psi})=\sum_{i=1}^{N}\sum_{(t,\omega)_{j}\in\Psi_{i}}\mathbb{E}_{\boldsymbol{\theta}\sim q_{\boldsymbol{\theta}}^{(k-1)}}\left[f(\omega_{j},-g^{\text{lin}}(t_{j},\mathbf{x}_{i};\boldsymbol{\theta}))\right]\\ +\sum_{i=1}^{N}\sum_{(t,\omega)_{j}\in\Psi_{i}}\mathbb{E}_{\phi\sim q_{\phi}^{(k-1)}}\left[\log\left(\frac{\phi}{Z(t_{j},\mathbf{x}_{i})}\right)\right]+\text{const.}$ (A20)
Plugging in the definition of $f(\cdot,\cdot)$ from (8) simplifies (A20) to
@@
$h^{\mathbb{Q},(k)}_{i}(t,\omega)$ $=\exp\left(-\frac{\left(\tilde{s}^{(k-1)}_{i}(t)\right)^{2}}{2}\omega\right)\,\cosh\left(\frac{\tilde{s}^{(k-1)}_{i}(t)}{2}\right),$
$\gamma_{i}^{\mathbb{Q},(k)}(t)$ $=\frac{1}{Z(t,\mathbf{x}_{i})}\sigma(\tilde{s}^{(k-1)}_{i}(t))\exp\left(-\frac{\tilde{m}^{(k-1)}_{i}(t)+\tilde{s}^{(k-1)}_{i}(t)}{2}+\mathbb{E}_{\phi\sim q^{(k-1)}_{\phi}}[\log\phi]\right),$
$\lambda^{\mathbb{Q},(k)}_{i}(t,\omega)$ $=\gamma_{i}^{\mathbb{Q},(k)}(t)h^{\mathbb{Q},(k)}_{i}(t,\omega)\lambda_{*,i}(t,\omega),$
where $\lambda_{*,i}(t,\omega)$ is the intensity defined in (A8). Furthermore, we define for convenience,
$\lambda_{i}^{\mathbb{Q},(k)}(t):=t^{\rho-1}\gamma_{i}^{\mathbb{Q},(k)}(t).$ (A21)
Notice that by using expression (A2), the function $\lambda^{\mathbb{Q},(k)}_{i}(t,\omega)$ can be written as
$\lambda^{\mathbb{Q},(k)}_{i}(t,\omega)=\lambda^{\mathbb{Q},(k)}_{i}(t)\>p_{\text{PG}}\left(\omega\mid 1,\tilde{s}^{(k-1)}_{i}(t)\right).$ (A22)
@@
Notice that the products $\gamma_{i}^{\mathbb{Q},(k)}(t_{j})h^{\mathbb{Q},(k)}_{i}(t_{j},\omega_{j})$ are all strictly positive  [^1] See Lemma N.1 for a proof of the strict positivity of the normalization factor $Z(t,\mathbf{x}_{i})$., hence $\frac{\mathrm{d}\mathbb{Q}_{\boldsymbol{\Psi}}^{(k)}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}$ is also strictly positive. Under suitable regularity conditions on $g$, one can show that $h^{\mathbb{Q},(k)}_{i}(t,\omega)$ and $\gamma_{i}^{\mathbb{Q},(k)}(t)$ satisfy the integrability criteria of (A9), so that $\mathbb{Q}_{\boldsymbol{\Psi}}^{(k)}$ is the probability measure under which each $\Psi_{i}$ $(i=1,\ldots,N)$ is a marked Poisson Process on $0,y_{i}]\times\mathbb{R}_{+}$ with intensity function $\lambda_{i}^{\mathbb{Q},(k)}(t,\omega)$.
@@ G.3 Optimal Update for $\phi$
Using standard mean-field variational inference techniques (see, e.g., Chapter 10.1 of [bishop_pattern_recognition]), the optimal variational factor for the parameter $\phi$ is obtained by computing the expectation of the joint log-density with respect to the other variational factors. In particular, we have
@@
Next, by substituting the expression for the augmented likelihood from (14), the Radon–Nikodym derivative of $\mathbb{P}_{\boldsymbol{\Psi}|\phi}$ with respect to $\mathbb{P}_{\boldsymbol{\Psi},*}$ from (A10), and the prior of $\phi$ from (2), we obtain,
@@
where $\lambda^{\mathbb{Q},(k)}_{i}(t)$ is shown in (A21). We deduce that
[…]
@@
$\tilde{\alpha}^{(k)}=\alpha_{0}+\sum_{i=1}^{N}\left(\delta_{i}+\int_{0}^{y_{i}}\lambda^{\mathbb{Q},(k)}_{i}(t)\mathrm{d}t\right),\quad\tilde{\beta}=\beta_{0}+\sum_{i=1}^{N}\int_{0}^{y_{i}}\frac{t^{\rho-1}}{Z(t,\mathbf{x}_{i})}\mathrm{d}t.$ (A23)
@@
[…]